OpenStudy (anonymous):

PLEASE HELP ME!!!! :(....Use the substitution method to solve the system of equations. 2x - 7y = 1 x - 8y = -4 A- (2, 4) B- (4, 2) C- (1, 4) D- (4, 1)

OpenStudy (anonymous):

so with the second equation x=-4+8y now sub that into first equation

OpenStudy (anonymous):

Huh?

OpenStudy (anonymous):

sub x=-4+8y into the first equation for where x is in the equation

OpenStudy (jonnyvonny):

With the substitution method, you want to get 1 variable ( a "x" or "y" [or any letter]) alone, then put that in in place of the other function which doesn't have that alone ie \[10x -2y = 6\]\[y=5x -3\] You would put the function "y=5x-3" into the other function, specifically where the "2y" is, since they're EQUAL. Now, with your problem, less steps would be required to get the "x" alone:\[x-8y=-4\rightarrow x=-4+8y\] Now that we have 1 variable alone, plug that variable, or letter, with the same letter in the other equation:\[x= -4+8y \left| \right| 2x - 7y = 1\rightarrow 2(-4+8y)-7y=1\rightarrow -8+16y-7y=1\] Now we simplify and get y-alone. \[-8+16y-7y=1\rightarrow -8+9y=1\rightarrow 9y=9\rightarrow y=1\] Since a point requires and x and y value, we may merely plug in the new-found y-value in any of the 2 original equations. Lets just do the first one becausethisisamerica \[y=1 \left| \right|2x-7y=1\rightarrow 2x-7(1)=1\rightarrow2x-7=1\rightarrow 2x=8\rightarrow x=4\] There we finallythefreakinggo :). The point of intersection, or the answer, is (4,1). Hope this helped, if you need more assistance, please feel free to pst me.